OpenStudy (anonymous):

find the equation of the tangent line to the graph of y=sqrt x/x+1 at point where x=1

OpenStudy (anonymous):

Step 1.)Plug in x =1 in the equation and find the y-coordinate. Step 2.)Take the derivative and evaluate it at x = 1. This will be your m in the next step. Once you have these two things, use y=mx+b or \[y-y _{1}=m(x-x _{1})\] to find the equation of the tangent line.

OpenStudy (anonymous):

im having a tough time finding the derivative for the equation, any way you could help?

OpenStudy (anonymous):

Could you re-write the equation properly? :D

OpenStudy (anonymous):

\[y=\sqrt{x}\div x+1\]

OpenStudy (anonymous):

\[y=\sqrt{x}/(x+1)\] is that the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This will be the derivative. \[(1-x)/(2(\sqrt(x)(x+1)^2)\]